# Hyperparameter Optimization in Machine Learning

### v2

@@ 1.6 Related surveys
The topic of this monograph has been surveyed in some other papers, briefly in [hutter_beyond_2015] and more extensively in  [bischl:2023:hyperparameter_optimization_foundations, yang:2020:hyperparameter_optimization_machine]. Arlot and Celisse arlot:2010:survey_crossvalidation_procedures surveyed the model selection problem from a complementary statistical perspective. Specific aspects have been reviewed in other review papers, in particular we would like to mention Bayesian optimization [Frazier:08, shahriari:2016:taking_human_out, wang:2023:recent_advances_bayesian], neural architecture search [ren:2021:comprehensive_survey_neurala, heuillet:2024:efficient_automation_neural], evalutionary algorithms [li:2023:survey_evolutionary_deep], and meta-learning [hospedales2020meta] HPO is a core topic in automated machine learning (AutoML). The subject is covered extensively in the book by Hutter et al. hutter2019automated and more recently in [he:2021:automl]
@@ 1.7 Notation
As usual, we denote by $\mathbb{R}$ and $\mathbb{N}$ the set of real and natural numbers, respectively. For $K\in\mathbb{N}^{+}$, we indicate by $[K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We reserve the Greek letter $\lambda$ to indicate hyperparameters and $f$ for the response function. We use $\mathbb{P}$ to indicate probability functions. Whenever confusion may arise, we further add a subscript like $\mathbb{P}_{x\sim\nu}$ to specify that the random variable $x$ evaluated by $\mathbb{P}$ follow the distribution $\nu$. We use $\mathbb{E}$ with the same convention to indicate the expected value. We mostly reserve the letter $p$ to indicate density or probability mass functions. We denote by $\mathcal{N}(\mu,\Sigma)$ the Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
@@
Here, we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We discuss the nested structure that characterizes an HPO problem and the archetypal settings encountered in machine learning, contrasting this to more general algorithm configuration problems. To fix ideas, we revisit in greater detail the example of the introduction and showcase different application scenarios from the literature. We conclude the chapter with an overview of practical desiderata for HPO algorithms that will guide our discussions in the remainder of the monograph.
@@ 2.1 Learning algorithms with hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm relies on in order to generalize outside of the training data. Formally, let:
  • $\mathscr{D}$ be a the collection of all finite datasets that the algorithm can process;
  • $\Lambda$ be an hyperparameter space, often though of as the Cartesian product of individual hyperparameters spaces;
  • $\mathcal{H}$ be a model space, typically indexed by a parameter vector $w$, that represents the class of models that the algorithm may output.
We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h_{\lambda}$, capable of accomplishing a given task:
$A:\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda).$ (2.1)
We invariably call a $\lambda\in\Lambda$ a hyperparameter vector or a hyperparameter configuration. By writing $h_{\lambda}=A(\mathcal{D},\lambda)$ in Eq. (2.1) we emphasize the dependence of the model returned by the algorithm on the hyperparameter vector $\lambda$, but omit the dependence on the training data $\mathcal{D}$. This choice is consistent with the fact that often one considers the dataset fixed and “given” when setting up an HPO problem.
@@ 2.1.1 Logistic regression for sentiment analysis: an HPO perspective
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Let us denote the data by $\mathscr{D}:=\bigcup_{N\in\mathbb{N}^{+}}(\mathcal{X}\times\mathcal{Y})^{N}$, where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ is the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda_{1}\times\dots\times\Lambda_{7}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}:\mathcal{X}\to\mathbb{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
@@
=$\mathcal{H}=\left\{h_{\lambda}:\mathcal{X}\to\mathcal{Y}\,:\,h_{\lambda}(x)=\textbf{1}_{\mathbb{R}^{+}}(w^{\intercal}\Phi_{\lambda}(x)),\;w\in\mathbb{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
@@
where $\mathbf{1}_{S}$ is the indicator function: $\mathbf{1}_{S}(y)=1$ if $y\in S$ and 00 otherwise. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ vary in their respective spaces $\Lambda_{1}$, $\Lambda_{2}$, $\Lambda_{3}$ and $\Lambda_{4}$.
@@
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the $n$-grams considered, the more input features and, consequently, parameters the model will have. This range is controlled by the hyperparameters $\lambda_{1}$ and $\lambda_{2}$. Unigrams alone ($\lambda_{1}=\lambda_{2}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer $n$-grams typically requires increasingly larger corpora to avoid overfitting, since the number of possible features (sequences of words of length $n$) grows exponentially with $n$.
@@
The remaining hyperparameters, $\lambda_{5},\lambda_{6}$ and $\lambda_{7}$, do not directly affect the model space, but rather determine the way in which the optimization is carried out. Specifically, the algorithm $A$ outputs the model whose parameter vector $w$ minimizes the training objective $\mathcal{J}_{\lambda}:\mathbb{R}^{n_{\lambda}}\times\mathscr{D}\to\mathbb{R}$ defined as follows:
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log\mathbb{P}\left(y|\Phi_{\lambda}(x);w\right)+\lambda_{6}\Omega_{\lambda_{5}}(w),$ (2.3)
@@
where $\Omega_{\lambda_{5}}(w)$ is either the $L^{1}$ or the $L^{2}$ norm depending on the value of $\lambda_{5}$, and the probability that the model assigns a document to express a positive sentiment, encoded by $y=1$, is given by
@@
Hyperparameter $\lambda_{5}$ exposes a binary choices between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values. Figure 2.1 offers a geometric interpretation of these regularization schemes. The hyperparameter $\lambda_{6}$ controls the strength of such regularization; if, for instance, $\lambda_{5}=L^{1}$, one can expect the number of 00’s in the solution vector to be proportional to the value of $\lambda_{6}$.
Figure 2.1: Geometrical representations of the $L^{2}$ (left) and $L^{1}$ (right) regularizers. $w^{*}_{0}$ represents the minimizer of the training loss $\mathcal{J}_{\lambda}$ when the regularization coefficient $\lambda_{6}$ is 00 (i.e. no regularization). $w^{*}_{\rho_{1}}$ and $w^{*}_{\rho_{2}}$ are, instead, the minimizers of the regularized losses for two different values of $\lambda_{6}=\rho_{i}$ for $i\in\{1,2\}$, with $0<\rho_{1}<\rho_{2}$. These corresponds to the projections of $w^{*}_{0}$ onto $L^{2}$ and $L^{1}$ balls with varying radii that are inversely proportional to $\lambda_{6}$.
Finally, $\lambda_{7}$ controls the convergence tolerance of the numerical solver used to minimize the objective (2.3).
@@
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda_{6}$ in our example could in principle take any non-negative real value: it does not need to be limited to the interval $[10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in (2.3) instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
@@ 2.2 The hyperparameter optimization problem
Hyperparameter configurations need to be evaluated according to a given performance metric, that we denote by $\mathscr{M}$. This metric takes as input a trained model and a validation set, and returns a scalar value, such as the validation error.  [^1] In Section 8.1 we discuss the case where $\mathscr{M}$ returns multiple scalars. The objective of an HPO problem, the response function, is then constructed from the performance metric. Specifically, given a learning algorithm $A:\mathscr{D}\times\Lambda\to\mathcal{H}$, a metric :×$\mathscr{M}:\mathcal{H}\times\mathscr{D}\to\mathds{R}$, a training set $\mathcal{D}$ and a validation set $\mathcal{V}$, the hyperparameter optimization problem is defined as follows:
@@
such $\mbox{such that}\quad h_{\lambda}$ $=$ $A(\mathcal{D},\lambda).$ (2.5)
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until convergence. Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of out-of-the-box continuous algorithms for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 where we discuss gradient-based HPO techniques that explicitly leverage this structure. A key observation is that we cannot collapse the optimization problem into one level, where we jointly minimize both model parameters and hyperparameters. For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda_{6}=10^{-5}$ and $\lambda_{7}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.

### v3

@@ 1.6 Related Surveys
The topic of this monograph has been surveyed in some other papers, briefly by hutter_beyond_2015 and more extensively in the works of bischl:2023:hyperparameter_optimization_foundations and yang:2020:hyperparameter_optimization_machine. arlot:2010:survey_crossvalidation_procedures surveyed the model selection problem from a complementary statistical perspective. Specific aspects have been reviewed in other papers, in particular Bayesian optimization by Frazier:08, shahriari:2016:taking_human_out, wang:2023:recent_advances_bayesian, neural architecture search by ren:2021:comprehensive_survey_neurala, heuillet:2024:efficient_automation_neural, evolutionary algorithms by li:2023:survey_evolutionary_deep, and meta-learning by hospedales2020meta. HPO is a core topic in automated machine learning (AutoML). The subject has been covered extensively in the book by hutter2019automated and more recently in the survey by he:2021:automl.
This monograph aims to provide a broad methodological introduction to the problem settings and principal algorithmic approaches in HPO. It is intended to support both practitioners aiming to understand the principles behind the methods they use or consider adopting, and researchers entering the field or seeking a reference on specific algorithm classes or aspects of hyperparameter optimization.
@@ 1.7 Notation
We denote by $\mathds{R}$ and $\mathds{N}$ the set of real and natural numbers, respectively. For $K$ a positive integer, we indicate by $K]=\{1,\dots,K\}$ the subset of the natural numbers from $1$ to $K$, included. We use calligraphic letters, like $\mathcal{D}$, to indicate sets, and script-style calligraphic, like $\mathscr{D}$ to indicate collections. We reserve the Greek letter $\lambda$ to denote hyperparameters, $\Lambda$ to indicate hyperparameter spaces, and $f$ to denote the response function. We use $\mathbb{P}$ to indicate a probability measure. Whenever confusion may arise, we further add a subscript like $\mathbb{P}_{x\sim\nu}$ to specify that the random variable $x$ evaluated by $\mathbb{P}$ follow the distribution $\nu$. We use $\mathbb{E}$ with the same convention to indicate the expected value. We mostly reserve the letter $p$ to indicate density or probability mass functions, and denote by $\mathcal{P}(\mathcal{X})$ the space of densities over the set $\mathcal{X}$. We denote by $\mathcal{N}(\mu,\Sigma)$ the multivariate Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$.
@@
Here, we formalize hyperparameter optimization (HPO) along the concepts introduced in the previous chapter. We discuss the nested structure that characterizes an HPO problem and the archetypal settings encountered in supervised machine learning. To fix ideas, we revisit in greater detail the example of the introduction and showcase different application scenarios from the literature. We conclude the chapter with an overview of practical desiderata for HPO algorithms that will guide our discussions in the remainder of the monograph.
@@ 2.1 Learning Algorithms with Hyperparameters
Hyperparameters are variables that configure the behavior of a learning algorithm, shaping and modifying its inductive bias, i.e., the set of assumptions the algorithm implicitly relies on in order to generalize. We view a learning algorithm $A$ as a higher-order, possibly stochastic, function that maps data and hyperparameters to a model $h_{\lambda}$, capable of accomplishing a given task:
$A\mathrel{\mathop{\ordinarycolon}}\mathscr{D}\times\Lambda\to\mathcal{H};\quad h_{\lambda}=A(\mathcal{D},\lambda)$ (2.1)
where $\mathscr{D}$ denotes the collection of all finite datasets that the algorithm can process, $\Lambda$ the hyperparameter space (the Cartesian product of the domains of individual hyperparameters), $\mathcal{H}$ the model space that represents the class of models that the algorithm may output, and $\lambda\in\Lambda$ a hyperparameter vector (or tuple). By writing $h_{\lambda}=A(\mathcal{D},\lambda)$ in Eq. (2.1) we emphasize the dependence of the model returned by the algorithm on $\lambda$, but omit the dependence on the training data $\mathcal{D}$. This choice is consistent with the fact that one often considers the dataset “fixed and given” when setting up an HPO problem.
@@ 2.1.1 Logistic Regression for Sentiment Analysis: An HPO Perspective
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Data in this case is a set of pairs $\{(x,y)\mid x\in\mathcal{X},y\in\mathcal{Y}\}$ where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ is the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda^{(1)}\times\dots\times\Lambda^{(7)}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathds{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
@@
$\mathcal{H}=\left\{h_{\lambda}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{Y}\mid h_{\lambda}(x)=\mathds{1}\!\left\{w^{\intercal}\Phi_{\lambda}(x)>0\right\},\;w\in\mathds{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
@@
where $\mathds{1}\!\left\{s\right\}$ is the indicator function: $\mathds{1}\!\left\{s\right\}=1$ if $s$ is true and 0 otherwise. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda^{(1)}$, $\lambda^{(2)}$, $\lambda^{(3)}$, and $\lambda^{(4)}$ vary in their respective domains $\Lambda^{(1)}$, $\Lambda^{(2)}$, $\Lambda^{(3)}$ and $\Lambda^{(4)}$.
@@
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the $n$-grams considered, the more input features and, consequently, parameters the model will have. This range is controlled by the hyperparameters $\lambda^{(1)}$ and $\lambda^{(2)}$. Unigrams alone ($\lambda^{(1)}=\lambda^{(2)}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer $n$-grams typically requires increasingly larger corpora to avoid overfitting, since the number of possible features (sequences of words of length $n$) grows exponentially with $n$.
@@
The remaining hyperparameters, $\lambda^{(5)},\lambda^{(6)}$, and $\lambda^{(7)}$, do not directly affect the model space, but rather determine the way in which the optimization is carried out. Specifically, the algorithm $A$ outputs the model whose parameter vector $w$ minimizes the training objective $\mathcal{J}_{\lambda}\mathrel{\mathop{\ordinarycolon}}\mathds{R}^{n_{\lambda}}\times\mathscr{D}\to\mathds{R}$ defined as follows:
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log\mathbb{P}\left(y|\Phi_{\lambda}(x);w\right)+\lambda^{(6)}\Omega_{\lambda^{(5)}}(w),$ (2.3)
@@
where $\Omega_{\lambda^{(5)}}(w)$ is either the $L^{1}$ or the $L^{2}$ norm depending on the value of $\lambda^{(5)}$, and the probability that the model assigns a document to express a positive sentiment, encoded by $y=1$, is given by
@@
Hyperparameter $\lambda^{(5)}$ exposes a binary choice between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values. Figure 2.1 offers a geometric interpretation of these regularization schemes. The hyperparameter $\lambda^{(6)}$ controls the strength of such regularization; if, for instance, $\lambda^{(5)}=L^{1}$, one can expect the number of 0’s in the solution vector to be proportional to the value of $\lambda^{(6)}$.
Figure 2.1: Geometrical representations of the $L^{2}$ (left) and $L^{1}$ (right) regularizers. $w^{*}_{0}$ represents the minimizer of the training loss $\mathcal{J}_{\lambda}$ when the regularization coefficient $\lambda^{(6)}$ is 0 (i.e. no regularization). $w^{*}_{\rho_{1}}$ and $w^{*}_{\rho_{2}}$ are, instead, the minimizers of the regularized losses for two different values of $\lambda^{(6)}=\rho_{i}$ for $i\in\{1,2\}$, with $0<\rho_{1}<\rho_{2}$. These correspond to the projections of $w^{*}_{0}$ onto $L^{2}$ and $L^{1}$ balls with varying radii that are inversely proportional to $\lambda^{(6)}$.
Finally, $\lambda^{(7)}$ controls the convergence tolerance of the numerical solver used to minimize the objective in Eq. (2.3).
@@
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda^{(6)}$ in our example could in principle take any non-negative real value: it does not need to be limited to the interval $10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in Eq. (2.3), instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
@@ 2.2 The Hyperparameter Optimization Problem
Hyperparameter configurations need to be evaluated according to a given performance metric, that we denote by $\mathscr{M}$. This metric takes as input a trained model and a validation set, and returns a scalar value, such as the validation error [^1] In Section 8.1 we discuss the case where $\mathscr{M}$ returns multiple scalars.. The objective of an HPO problem, the response function, is then constructed from the performance metric. Specifically, given a learning algorithm $A\mathrel{\mathop{\ordinarycolon}}\mathscr{D}\times\Lambda\to\mathcal{H}$, a metric $\mathscr{M}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\times\mathscr{D}\to\mathds{R}$, a training set $\mathcal{D}$ and a validation set $\mathcal{V}$, the hyperparameter optimization problem is defined as follows:
@@
$\mbox{such that}\quad h_{\lambda}$ $=$ $A(\mathcal{D},\lambda).$ (2.5)
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until termination (e.g. convergence). Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of out-of-the-box continuous algorithms for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, Problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 where we discuss gradient-based HPO techniques that explicitly leverage this structure. A key observation is that we cannot collapse the optimization problem into one level, where we jointly minimize both model parameters and hyperparameters. For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda^{(6)}=10^{-5}$ and $\lambda^{(7)}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.